Find the third degree Taylor Polynomial for the function f(x) = cos x at a = −π/4.
Answer #1

We can use Taylor series formula

T_3(x)=f(a)+(x-a)f'(a)+\frac{(x-a)^2}{2!}f''(a)+\frac{(x-a)^3}{3!}f'''(a)

a = −π/4.

f(x) = cos x

now, we can find values and plug in formula

f(-\frac{\pi}{4})=\frac{\sqrt{2}}{2}

f'(x)=-sin(x)

f'(-\frac{\pi}{4})=-sin(\frac{-\pi}{4})=\frac{\sqrt{2}}{2}

f''(x)=-cos(x)

f''(-\frac{\pi}{4})=-cos(-\frac{\pi}{4})=-\frac{\sqrt{2}}{2}

f'''(x)=sin(x)

f'''(-\frac{\pi}{4})=sin(-\frac{\pi}{4})=-\frac{\sqrt{2}}{2}

now, we can plug values

and we get

T_3(x)=\frac{\sqrt{2}}{2}+\frac{1}{\sqrt{2}}\left(x+\frac{\pi }{4}\right)-\frac{1}{2\sqrt{2}}\left(x+\frac{\pi }{4}\right)^2-\frac{1}{6\sqrt{2}}\left(x+\frac{\pi }{4}\right)^3...........Answer